Let’s say we replace a network of capacitors with one capacitor of value Ceq. In terms of charge, voltage, and energy, can Ceq truly represent the entire network?  Explain why or why not.
Answer #1

Yes the equivalent capacitor Ceq can represent the entire network, bacause charge Qeq= Q1+Q2+............; Veq= V1+V2+..............; total energy will also be conserved because Eeq=1/2*Ceq*Veq2

answered by: ANURANJAN SARSAM